OpenStudy (darkbluechocobo):

A longitudinal wave has a period of 10 seconds. What is the frequency

OpenStudy (darkbluechocobo):

@Data_LG2

OpenStudy (anonymous):

Period is inversely proportional to frequency. The formula is T=1/f, hence 10=1/f and making as subject becomes f=1/10 or 0.1Hz.

OpenStudy (darkbluechocobo):

Ugh there are somany formulas to memorize T_T

OpenStudy (anonymous):

About waves? Just prctise with the book or exam paper questions related to such calculation, so that you get used to them.